OpenStudy (anonymous):

A population of rhinos decreases by 22% each year how many rhino's will there be in 4 years if there are 200 rhinos now ?

OpenStudy (anonymous):

200 - 22% = X. X - 22% =... It's pretty simple.

OpenStudy (anonymous):

is x a number ?

OpenStudy (anonymous):

200 - 22% = 156 then you subract 22% of 156

OpenStudy (anonymous):

okay thanks :)

OpenStudy (anonymous):

If you want I can finish the problem and just give you the answer.

OpenStudy (anonymous):

that sounds good

OpenStudy (anonymous):

The answer is\[ 200 (.78)^4\approx 74 \]